OpenStudy (anonymous):

what is the radian measure of the angle formed by the hands of a clock at 2

OpenStudy (anonymous):

i got it pi over three right

OpenStudy (e.cociuba):

When the long hand of the clock makes 1 revolution, it has gone from12 to 12. One revolution is divided into 12 parts. 2 parts out of 12 = 2/12 = 1/6 of the circumference. One circumference = 2π radians. What is the radian measure of the angle formed by the hands of a clock at 2:00 p.m.? 2/6 * (2π radians) 1/6 radian

OpenStudy (anonymous):

tell me if this makes sense

OpenStudy (anonymous):

3 oclock is 90 degrees

OpenStudy (e.cociuba):

k wat??

OpenStudy (anonymous):

2/3 of 90 degrees is 60 degrees

OpenStudy (anonymous):

60 degrees converted to radians is pi/3 correct

OpenStudy (anonymous):

?

OpenStudy (anonymous):

its in terms of pi

OpenStudy (e.cociuba):

Yes tht is right. It makes sense

OpenStudy (anonymous):

thanQ

OpenStudy (e.cociuba):

np:)